OpenStudy (icalibear):

Did I do this right? Please check my work, somebody? On January 1, 2010, Chessville has a population of 50,000 people. Chessville then enter a period of population growth. Its population increases 7% each year. On the same day, Checkersville has a population of 70,000 people. Checkersville starts to experience a population decline. Its population decreases 4% each year. During what year will the population of Chessville first exceed that of Checkersville? Please show your work and explain your steps in the simplest way possible. :)

OpenStudy (anonymous):

(50K)(.07)=3500 (70K)(.04)=2800 in y1 of you enter 50000 + 3500x and y2 you enter 70000 - 2800x...go to the table and you will see that they equal after about 3 years

OpenStudy (anonymous):

first you want to convert the percentages to decimals

OpenStudy (anonymous):

so you divide 7 by 100 and 4 by 100

OpenStudy (anonymous):

7/100= .07 4/100= .04

OpenStudy (anonymous):

then you want to multiply 50,000 by .07 and 70,000 by .04

OpenStudy (anonymous):

50,000 times .07=3500 70,000 times .04= 2800

OpenStudy (anonymous):

50,000 times .07=3500 70,000 times .04= 2800

OpenStudy (icalibear):

thanks